OpenStudy (anonymous):

A fruit-seller purchased 720 oranges at the rate of Rs 6.50 per dozen. 30 of them were found to be damaged. At what price should he sell the remaining oranges so as to get overall profit of 15%.

OpenStudy (anonymous):

plz tell fast.

OpenStudy (anonymous):

i m late.. g2g tomorrow is school bye, thnx for trying!

OpenStudy (anonymous):

720 oranges= 60 dozens so total rate= 60*6.5rupees=.... now 720 are at a price of ..... now you have actual 690 oranges at cost price of .... so you have cost price * (1+15/100)=selling price now devide selling price by 690 to have selling price of 1 orange

OpenStudy (anonymous):

@kajolc.2496

OpenStudy (across):

OpenStudy is not a homework-answering machine.

OpenStudy (anonymous):

its not my homework!

OpenStudy (anonymous):

get it!!

OpenStudy (across):

You're right: it's ours.

OpenStudy (anonymous):

lol @across leave it she is just a kid. . . i am explaining formula and let her do the rest . . .

OpenStudy (anonymous):

its my brother's homework, if i'll not do he'll kill me!

OpenStudy (anonymous):

lol i can understand. . . go now

OpenStudy (anonymous):

i m jst a 10 yr girl

OpenStudy (anonymous):

ale bacha chalo jao ab school jana hoga subah 5th me ho??

OpenStudy (anonymous):

will anyone plz tell me the answer! its urgent plz...

OpenStudy (anonymous):

@kajolc.2496...i think u shud give up nd @theyatin....i consider that trolling.